Consider the following reaction, which occurs at standard state: 2Cl(g) --> Cl2(g) ∆H° = -- -243.4 kJ, if 15.0g of Cl is considered to chlorine, determine the heat released from this reaction. I need this urgently please!


1
Expert's answer
2021-02-03T07:22:40-0500

n(Cl) = 15/35.5 = 0.4225 mol

2 mol Cl ------- -243.4 kJ

0.4225 mol Cl ------ ∆H

∆H = 0.4225*(-243.4)/2 = -51.418 kJ

Answer: -51.418 kJ